OpenStudy (anonymous):

**WILL MEDAL** Complete the standard form of the equation -4/3y=3-x ?x +?y = 9

OpenStudy (anonymous):

is it \[-\frac{4}{3}y=3-x\]?

OpenStudy (haleyelizabeth2017):

I think it is.......so you all you need to do is move x to the left of the equal sign....do you know how to do that?

OpenStudy (anonymous):

i might multiply both sides by 3 first and get \[-4y=9-3x\] then add the \(3x\) to both sides to finish it

OpenStudy (haleyelizabeth2017):

wait......let me see....:/ it looks like they got rid of the fraction by multiplying everything by a certain number, which you will need to figure out :D

OpenStudy (anonymous):

Sorry my computer froze

OpenStudy (anonymous):

Is it 9x+-7x=9?

OpenStudy (anonymous):

@haleyelizabeth2017